# Improving Sequential Recommendations with LLMs

### v1

@@ 1. Introduction
Lately, the advent of Large Language Models (LLM), which in an auto-regressive mode repeatedly complete a sequence until they fulfil a prompt task (Radford et al., 2018), has opened a fruitful research direction for sequential recommendation. In fact, LLMs based on Generative Pretrained Transformers (Radford et al., 2019) are only the latest of a stream of innovations in Natural Language Processing (NLP) that inspired the birth of new sequential recommendation models. Notable examples are GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019), which were influenced by the Gated Recurrent Unit (GRU) (Cho et al., 2014), the transformer architecture (Vaswani et al., 2017), and BERT (Devlin et al., 2019), respectively. The key contribution of LLMs to the sequential recommendation problem is a semantic model of real world concepts and their representation in natural language sequences. This consideration is neglected by typical sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019).
The rapidly growing interest for involving LLMs in recommender systems clusters around cross-domain (Zhang et al., 2021a; Hou et al., 2023b), domain-specific (Bao et al., 2023; Zhang and Wang, 2023), and multi-modal (Wu et al., 2022; Wang et al., 2022) recommendations. Within the scope of recommendation approaches, LLMs are involved either by providing item embeddings (Hou et al., 2023a) or by responding to a recommendation task conveyed in a prompt, which may contain zero (Ding et al., 2022) or a few examples (Geng et al., 2022) to enrich the context of the task available to an LLM. Alternatively, an LLM can be fine-tuned to learn how to perform better for a specific recommendation task (Hou et al., 2022). Regarding the nature of recommendation tasks posed to LLMs, they are typically used to generate (Li et al., 2023f) or rerank (Hou et al., 2023b) recommendations. Altogether, the current state of research is summarized in the following surveys (Lin et al., 2023; Wu et al., 2023). Given the currently early stage of this research avenue, the prospect for the various recommendation approaches involving LLMs remains largely unexplored.
In this work, [^2] The present paper substantially extends our previous work on leveraging LLMs for the sequential recommendation problem (Harte et al., 2023). We detail our introduced extensions in Section 2. we design and research the potential of three orthogonal approaches and hybrids of those that leverage LLMs for sequential recommendations. We decompose each approach into its constituent technical aspects, identify or devise alternatives for each aspect, and investigate the impact of choices to the approach’s performance. Finally, we thoroughly tune and evaluate all approaches under the same testbed across three datasets of different domains and characteristics. These datasets are Amazon Beauty from the e-commerce beauty domain, a proprietary Delivery Hero dataset from the e-commerce darkstore domain, and Steam from the gaming domain.
@@
Our work results in the following contributions and insights.
  • We design and research three orthogonal methods and two hybrid approaches of leveraging LLMs for sequential recommendation. In particular, we investigate the technical aspects of each method, determine alternatives for each, thoroughly tune them, and evaluate their impact.
@@ 2. Background & Related Work
In this work we explore the potential of using LLMs for sequential recommendation problems (Jannach et al., 2020). In these problems, we consider as input a sequence of user interactions (i.e.a session) $S^{u}=(S^{u}_{1},S^{u}_{2},\ldots,S^{u}_{n})$, where $u$ is a user, $n$ is the length of the sequence and $S^{u}_{i}$ are individual items. The aim is to predict the next interaction of the given sequence. Besides the recent sequential recommendation models mentioned in the introduction (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019), in earlier works the sequential recommendation problem has been modelled as a Markov Chain (Garcin et al., 2013) or a Markov Decision Process (Shani et al., 2005). Neighborhood-based approaches, such as SKNN (Jannach and Ludewig, 2017), have also been proposed.
@@ 2.2. Domain-specific LLM-based Sequential Recommendation Models
In the movie, book, and e-commerce domains, the lightweight TALLRec framework proposed by Bao et al(Bao et al., 2023) aims to adapt LLMs for recommendations by structuring the recommendations as instructions, and tuning the LLMs through an instruction-tuning process. Experiments using the LLaMA model (Touvron et al., 2023) demonstrate that TALLRec significantly advances the LLM capabilities, and is highly efficient even on a single conventional GPU. Similarly, PALR (Yang et al., 2023) employs user/item interactions for candidate retrieval, and a fine-tuned LLM-based ranking model to produce recommendations. Using LLaMA for the LLM, and compared to representative baselines including GRU4Rec (Hidasi et al., 2016) and SASRec (Kang and McAuley, 2018), PALR appears more effective in various sequential recommendation tasks. Furthermore, GPTRec (Petrov and Macdonald, 2023), which is based on the GPT-2 architecture (Radford et al., 2019), employs a novel SVD tokenization algorithm and a Next-K recommendation strategy to generate sequential recommendations, accounting for already recommended items. GPTRec matches the quality of SASRec (Kang and McAuley, 2018) but is more GPU-memory efficient due to its sub-item tokenization. Inspired by search engines, GPT4Rec (Li et al., 2023f) generates a set of search queries based on the item titles from a user’s history, and then recommends items by searching these queries using beam search. Using GPT-2 (Radford et al., 2019) as a language model to generate queries that are used in a search engine using BM25 (Robertson and Zaragoza, 2009) as its score function, GPT4Rec outperforms state-of-the-art models including YouTubeDNN (Covington et al., 2016) and BERT4Rec (Sun et al., 2019). In addition, multi-query generation with beam search increases the diversity of recommended items and the coverage of a user’s interests.
@@ 3. LLMSeqSim: Semantic Item Recommendations via LLM Embeddings
We identified four technical aspects in this processing pipeline, where different implementation alternatives are possible. Below, in Sections 3.13.4, we elaborate our investigation of the different aspects involved in the LLMSeqSim approach.
@@ 4. LLMSeqPrompt: Prompt-based Recommendations by a Fine-Tuned LLM
In this approach, we inject domain knowledge to the collective information that a base LLM incorporates, with the goal of combining fundamental information about the items of the domain encoded in the LLM with domain-specific information conveyed in the fine-tuning process. The composition of these two elements of information exposed in the same highly capable substrate of an LLM paves the path towards potentially advanced recommendations.
To this end, we fine-tune an OpenAI ada [^7] https://platform.openai.com/docs/guides/fine-tuning model, an OpenAI GPT-3.5 turbo [^8] https://platform.openai.com/docs/guides/fine-tuning/what-models-can-be-fine-tuned model and the model text-bison@001999https://cloud.google.com/vertex-ai/docs/generative-ai/models/tune-models, which is based on Google PaLM2. GPT succeeded ada as OpenAI’s recommended base model for fine-tuning, while PaLM2 is a base model created by Google. While there is general consensus on the overall architecture and training process of a base LLM, differences in architectural choices, its configuration, and the training data mean that LLMs can differ significantly. [^10] A discussion of those factors is out of this paper’s scope. Thus, the aforementioned selection of models enables us to compare and contrast two fundamentally different LLMs and also assess the performance evolution and qualitative characteristics of OpenAI models, going from ada to GPT 3.5 turbo.
We fine-tune the models on training samples consisting of a prompt (the input) and a completion (the intended output). First, we take the training sessions and format them as prompt-completion pairs according to the fine-tuning task specification. We present the precise task specifications in Sections 4.14.24.3, and 4.4. We also separate a small subset of the training sessions and provide them to the fine-tuning process as validation set. We supply 500 sessions for validation to OpenAI and 250 sessions to Google PaLM, which is the maximum number of validation cases that the PaLM API allows. Then, we supply the prompt-completion pairs to a model’s fine-tuning API. In all configurations, we fine-tune the model until the validation loss converges in order to achieve the best performance.
@@
Besides employing a set of diverse LLMs, we propose four alternative specifications for the fine-tuning task. Our rationale is to examine the behavior and performance of a model on different problem formulations, namely generation of a single recommended item (Section 4.1), generation of a list of item recommendations (Section 4.2), classification (Section 4.3), and ranking (Section 4.4). [^12] Budget and time limitations allowed us to apply the fine-tuning tasks of Sections 4.24.3, and 4.4 to the Beauty dataset only. Generation, classification, and ranking enable an investigation of fine-tuning from less to more restricted settings. Fine-tuning a model on a generation task allows the model to recommend virtually anything. In a classification task, in contrast, a fine-tuned model is presented with a candidate pool of item recommendations presented as categories and selects a top-$k$ subset of those similarly to multi-class classification. Finally, the ranking task further narrows the freedom of the model in that the set of available recommendations is provided in the prompt, allowing the model to focus on the best ranking of those recommendations.
@@
4.1. Fine-tuning Task: Generate Single Recommended Item
In this variant, the prompt is a session, which contains a list of item metadata, e.g., product names, except for the last item, and the completion contains the metadata of the last item in the same session according to the first step in Figure 3. At prediction time, following the fine-tuning process, we repeat a prompt $k$ times ($k=20)$ and collect a single recommendation produced by the model per prompt invocation according to the fine-tune task formulation. We use the tendency of the model to make the same recommendation in different invocations for the same prompt as a proxy of its confidence for the recommendation, collect and deduplicate the recommendations over the $k$ invocations and rank the recommendations by frequency of appearance. The fine-tuned LLM, being a generative model, may also return hallucinated products. To remedy that, in the fifth step, we retrieve the embedding of each deduplicated recommendation using either the OpenAI or the Google embeddings API (this embedding might correspond to a hallucinated product), and finally we take the catalog’s product that is closest in terms of embedding similarity using the dot product.
@@ 4.1. Fine-tuning Task: Generate Single Recommended Item
4.2. Fine-tuning Task: Generate List of Item Recommendations
By design, a recommendation model aims to produce a number of desirable options in order to maximize the probability of satisfying a user’s intent at the time. Therefore, we would like to teach an LLM to embrace the role of a typical recommendation model and generate a list of next-item recommendations ranked by order of relevance that is determined by the fine-tuned model.
To convey this task in the fine-tuning process as we depict in Figure 4, we provide in the prompt a session containing a list of items except the last one. In the prompt completion we provide a list of item recommendations that is compiled by a state of the art recommendation model. The rationale is to ask a proven recommendation model to provide recommendations for the training sessions and use those as prompt completions for the fine-tune process. In fact, we can query more than one recommendation model and use the recommendations of the model that achieves the highest NDCG, or another metric, per training session. With enough training sessions, this practice could potentially transfer the knowledge of a powerful ensemble of models to a fine-tuned LLM. [^13] We did not apply this step in this work. Instead, we used LLMSeqSim in the Beauty dataset due to its high performance. Finally, we insert the ground truth item at the top of the list, if not already there, and fine-tune an LLM with the same prompts and the corresponding compiled list of recommendations as completion.
@@ 4.3. Fine-tuning Task: Classify Items for Next Item Recommendations
Asking an LLM to generate responses without explicitly providing the eligible options in the prompt represents the most challenging type of task. The reason is that it forces the model to choose from a virtually infinite set of options often leading to hallucinations. Therefore, we aim to liberate an LLM from the complexity of generating unbounded responses based on its fine-tuning. Instead, we allow it to focus on picking and ranking a list of pertinent item recommendations from an overall pool of items that we pose to the model as classification categories. The intuition behind this formulation is that there exists a modest subset of popular and diverse items of a domain that will please most customers avoiding, at the same time, niche items that receive scant attention.
In this variant, depicted in Figure 5, we state the next item recommendation problem as a multi-class classification task as follows. Prior to the fine-tuning process, we apply a clustering algorithm, namely K-means, to group items to clusters based on the embedding distances of their metadata [^14] We use product names in the Amazon Beauty dataset. As a result, the clusters can be seen as a set of diverse item groups that are potentially representative of the whole item space. We pick a relatively large number of clusters, i.e., 200, to maximize the probability of a good representation of semantically different items by the corresponding clusters. Then we elevate the most popular item of a cluster to represent the group of items in the cluster. In this fashion, we pick an array of popular and diverse candidate items that could represent good-quality recommendations across the total item space. Finally, for each training session, we take the embeddings of the ground truth item and the items that represent the classification categories and select the top-$k$ semantically closest classification categories for each ground truth item.
In the prompts used for fine-tuning, together with each input session of items (except the last item), we include the classification categories and in the corresponding completions we include the top-$k$ classification categories that we computed in the previous step. At prediction time, we ask the model to classify a prompt session to top-$k$ categories, which correspond to next item recommendations. Note that since the candidate recommendations are included in the prompt, there is no need for an extra step to map the recommendations back to the item space. The fine-tuned models respond according to their fine-tuning task specification.

### v2

@@ 1. Introduction
Lately, the advent of Large Language Models (LLM), which in an auto-regressive mode repeatedly complete a sequence until they fulfil a prompt task (Radford et al., 2018), has opened a fruitful research direction for sequential recommendation. In fact, LLMs based on Generative Pretrained Transformers (Radford et al., 2019) are only the latest of a stream of innovations in Natural Language Processing (NLP) that inspired the birth of new sequential recommendation models. Notable examples are GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019), which were influenced by the Gated Recurrent Unit (GRU) (Cho et al., 2014), the transformer architecture (Vaswani et al., 2017), and BERT (Devlin et al., 2019), respectively. The key contribution of LLMs to the sequential recommendation problem is a semantic model of real world concepts and their representation in natural language sequences. This consideration is neglected by typical ID-based sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019).
The rapidly growing interest for involving LLMs in recommender systems clusters around cross-domain (Zhang et al., 2021a; Hou et al., 2023b), domain-specific (Bao et al., 2023; Zhang and Wang, 2023), and multi-modal (Wu et al., 2022; Wang et al., 2022) recommendations. Within the scope of recommendation approaches, LLMs are involved either by providing item embeddings (Hou et al., 2023a) or by responding to a recommendation task conveyed in a prompt, which may contain zero (Ding et al., 2022) or a few examples (Geng et al., 2022) to enrich the context of the task available to an LLM. Alternatively, an LLM can be fine-tuned to learn how to perform better for a specific recommendation task (Hou et al., 2022). Regarding the nature of recommendation tasks posed to LLMs, they are typically used to generate (Li et al., 2023f) or rerank (Hou et al., 2023b) recommendations. Altogether, the current state of research is summarized in the following surveys (Lin et al., 2023; Wu et al., 2023).
In this paper, we build on the ideas developed in these early works on involving LLMs in recommender systems, and we design and research the potential of three orthogonal approaches and hybrids of those that leverage LLMs for the particular problem setting of sequential recommendations. We decompose each approach into its constituent technical aspects, identify or devise alternatives for each aspect, and investigate the impact of choices to the approach’s performance. Finally, we thoroughly tune and evaluate all approaches under the same testbed across three datasets of different domains and characteristics. These datasets are Amazon Beauty from the e-commerce beauty domain, a proprietary Delivery Hero dataset from the e-commerce darkstore domain, and Steam from the gaming domain.
@@
Our work results in the following contributions and insights. [^2] The present paper substantially extends our previous work on leveraging LLMs for the sequential recommendation problem (Harte et al., 2023). We detail our introduced extensions in Section 2.
  • Building on the growing literature regarding the use of LLMs in recommender systems, we design and research three orthogonal methods and two hybrid approaches of leveraging LLMs for sequential recommendation. In particular, we investigate the technical aspects of each method, determine alternatives for each, thoroughly tune them, and evaluate their impact.
@@
Overall, the main contribution of our work is an in-depth analysis of the potential of various previously-explored and novel ways of leveraging LLMs for sequential recommendation problems. Our comprehensive analyses, which involve different datasets, three LLMs, and a set of orthogonal technical approaches reveal that substantial performance improvements can indeed be obtained by involving LLMs for sequential recommendations.
@@ 2. Background & Related Work
In this work we explore the potential of using LLMs for sequential recommendation problems (Jannach et al., 2020). In these problems, we consider as input a sequence of user interactions (i.e.a session) $S^{u}=(S^{u}_{1},S^{u}_{2},\ldots,S^{u}_{n})$, where $u$ is a user, $n$ is the length of the sequence and $S^{u}_{i}$ are individual items. The aim is to predict the next interaction of the given sequence. Besides the recent sequential recommendation models mentioned in the introduction (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019), S3-Rec (Zhou et al., 2020) pre-trains a transformer architecture enhanced with self-supervised signals aiming to learn better data representations using the mutual information maximization principle. In earlier works the sequential recommendation problem has been modelled as a Markov Chain (Garcin et al., 2013) or a Markov Decision Process (Shani et al., 2005). Neighborhood-based approaches, such as SKNN (Jannach and Ludewig, 2017), have also been proposed.
@@ 2.2. Domain-specific LLM-based Sequential Recommendation Models
In the movie, book, and e-commerce domains, the lightweight TALLRec framework proposed by Bao et al(Bao et al., 2023) aims to adapt LLMs for recommendations by structuring the recommendations as instructions, and tuning the LLMs through an instruction-tuning process. Experiments using the LLaMA model (Touvron et al., 2023) demonstrate that TALLRec significantly advances the LLM capabilities, and is highly efficient even on a single conventional GPU. Similarly, PALR (Yang et al., 2023) employs user/item interactions for candidate retrieval, and a fine-tuned LLM-based ranking model to produce recommendations. Using LLaMA for the LLM, and compared to representative baselines including GRU4Rec (Hidasi et al., 2016) and SASRec (Kang and McAuley, 2018), PALR appears more effective in various sequential recommendation tasks.
Furthermore, GPTRec (Petrov and Macdonald, 2023), which is based on the GPT-2 architecture (Radford et al., 2019), employs a novel SVD tokenization algorithm and a Next-K recommendation strategy to generate sequential recommendations, accounting for already recommended items. GPTRec matches the quality of SASRec (Kang and McAuley, 2018) but is more GPU-memory efficient due to its sub-item tokenization. Inspired by search engines, GPT4Rec (Li et al., 2023f) generates a set of search queries based on the item titles from a user’s history, and then recommends items by searching these queries using beam search. Using GPT-2 (Radford et al., 2019) as a language model to generate queries that are used in a search engine using BM25 (Robertson and Zaragoza, 2009) as its score function, GPT4Rec outperforms state-of-the-art models including YouTubeDNN (Covington et al., 2016) and BERT4Rec (Sun et al., 2019). In addition, multi-query generation with beam search increases the diversity of recommended items and the coverage of a user’s interests.
@@ 3. LLMSeqSim: Semantic Item Recommendations via LLM Embeddings
Overall, LLMSeqSim shares similarities with the sequential model proposed in (Li et al., 2023c), which also relies on creating item embeddings from textual representations derived from metadata features, although not based on LLMs. We identified four technical aspects in this processing pipeline, where different implementation alternatives are possible. Below, in Sections 3.13.4, we elaborate our investigation of the different aspects involved in the LLMSeqSim approach.
@@ 4. LLMSeqPrompt: Prompt-based Recommendations by a Fine-Tuned LLM
In this approach, we inject domain knowledge to the collective information that a base LLM incorporates. Our goal is to thereby combine fundamental information about the items of the domain encoded in the LLM with domain-specific information conveyed in the fine-tuning process. Related approaches to fine-tune pre-trained LLMs for recommendation problems in different ways were successfully explored in (Bao et al., 2023; Zhang et al., 2023; Wang et al., 2022).
To this end, we fine-tune an OpenAI ada [^7] https://platform.openai.com/docs/guides/fine-tuning model, an OpenAI GPT-3.5 turbo [^8] https://platform.openai.com/docs/guides/fine-tuning/what-models-can-be-fine-tuned model, and the model text-bison@001999https://cloud.google.com/vertex-ai/docs/generative-ai/models/tune-models, which is based on Google PaLM2. GPT succeeded ada as OpenAI’s recommended base model for fine-tuning, while PaLM2 is a base model created by Google. While there is general consensus on the overall architecture and training process of a base LLM, differences in architectural choices, its configuration, and the training data mean that LLMs can differ significantly. Thus, the three selected models enable us to compare and contrast two fundamentally different LLMs, and to also assess the performance evolution of the two OpenAI models.
We fine-tune the models on training samples consisting of a prompt (the input) and a completion (the intended output). First, we take the training sessions and format them as prompt-completion pairs according to the fine-tuning task specification. [^10] We present the precise task specifications in Sections 4.1 to Section 4.4. We also separate a small subset of the training sessions and provide them to the fine-tuning process as validation set. We supply 500 sessions for validation to OpenAI and 250 sessions to Google PaLM, which is the maximum number of validation cases that the PaLM API allows. Then, we supply the prompt-completion pairs to a model’s fine-tuning API. In all configurations, we fine-tune the model until the validation loss converges in order to achieve the best performance.
@@
Besides employing a set of diverse LLMs, we propose four alternative specifications for the fine-tuning task. Our rationale is to examine the behavior and performance of a model on different problem formulations, namely generation of a single recommended item (Section 4.1), generation of a list of item recommendations (Section 4.2), classification (Section 4.3), and ranking (Section 4.4). [^12] Budget and time limitations allowed us to apply the fine-tuning tasks of Sections 4.24.3, and 4.4 to the Beauty dataset only. Generation, classification, and ranking enable an investigation of fine-tuning from less to more restricted settings. Fine-tuning a model on a generation task allows the model to recommend virtually anything. In a classification task, in contrast, a fine-tuned model is presented with a candidate pool of item recommendations presented as categories and selects a top-$k$ subset of those similarly to multi-class classification. Finally, the ranking task further narrows the freedom of the model in that the set of available recommendations is provided in the prompt. This allows the model to focus on the best ranking of those recommendations.
@@
4.1. Fine-tuning Task: Generate a Single Recommended Item
In this variant, the prompt is a session, which contains a list of item metadata, e.g., product names, except for the last item, and the completion contains the metadata of the last item in the same session according to the first step in Figure 3. At prediction time, following the fine-tuning process, we repeat a prompt $k$ times ($k=20)$ and collect a single recommendation produced by the model per prompt invocation according to the fine-tune task formulation. We use the tendency of the model to make the same recommendation in different invocations for the same prompt as a proxy of its confidence for the recommendation. We then collect and deduplicate the recommendations over the $k$ invocations and rank the recommendations by frequency of appearance. The fine-tuned LLM, being a generative model, may also return hallucinated products. To remedy that, in the fifth step, we retrieve the embedding of each deduplicated recommendation and finally take the catalog’s product that is closest in terms of embedding similarity using the dot product.
@@ 4.1. Fine-tuning Task: Generate a Single Recommended Item
4.2. Fine-tuning Task: Generate a List of Item Recommendations
In this approach, we teach an LLM to embrace the role of a typical recommendation model and generate a list of next-item recommendations ranked by order of relevance.
To convey this task in the fine-tuning process as we depict in Figure 4, we provide in the prompt a session containing a list of items except the last one. In the prompt completion we provide a list of item recommendations that is compiled by a state of the art recommendation model. The rationale is to ask a proven recommendation model to provide recommendations for the training sessions and use those as prompt completions for the fine-tuning process. In our case, we used LLMSeqSim in the Beauty dataset due to its high performance. [^13] In theory, we could also query more than one recommendation model and use the recommendations of the model that achieves the highest NDCG, or another metric. Finally, we insert the ground truth item at the top of the list, if not already there, and fine-tune an LLM with the same prompts and the corresponding compiled list of recommendations as completion.
@@ 4.3. Fine-tuning Task: Classify Items for Next Item Recommendations
In this approach, we explicitly provide the LLM with the eligible options for recommendation, instead of allowing it to ‘freely’ return recommendations. The learning goal is thus to select and rank a set of pertinent item recommendations from an overall pool of items that we pose to the model as classification categories. The intuition behind this formulation is that there exists a small subset of popular and diverse items in each domain that will please most customers. Restricting the set of options this way also helps us to avoid niche items that receive scant attention.
We state the next item recommendation problem as a multi-class classification task as follows, see also Figure 5. Prior to the fine-tuning process, we apply a clustering algorithm, namely K-means, to group items to clusters based on the embedding distances of their metadata. [^14] We use product names in the Amazon Beauty dataset. The resulting clusters can be seen as a set of diverse item groups that are representative of the whole item space. We pick a relatively large number of clusters, i.e., 200, to maximize the probability of a good representation of semantically different items. Then we elevate the most popular item of a cluster to represent the group of items in the cluster. In this fashion, we pick an array of popular and diverse candidate items that could represent good-quality recommendations across the total item space. Finally, for each training session, we take the embeddings of the ground truth item and the items that represent the classification categories and select the top-$k$ semantically closest classification categories for each ground truth item.
The prompts for fine tuning include the input session of items (except) the last item and the classification categories. In the corresponding completions, we include the top-$k$ classification categories that we computed in the previous step, see Figure 5 for an example. At prediction time, we ask the model to classify a prompt session to top-$k$ categories, which correspond to next item recommendations. Note that since the candidate recommendations are included in the prompt, there is no need for an extra step to map the recommendations back to the item space.